OpenStudy (anonymous):

Find the equations of thetangents to the follwoing graphs for the given values of x. (a) y=Inx (where x= 1/2) (b) y=In2x (where x=1/2)

OpenStudy (campbell_st):

differentiate and then plug in x = 1/2 y = ln x dy/dx = 1/x so 1/(1/2) = 2 then gradient of the tangent is 2 at x = 1/2 y = 2ln(2x) dy/dx = 2 x 2/2x = 2/x at x = 1/2 the gradient is 4

OpenStudy (anonymous):

2x-In2-1 is the answer for the first question... I did as you said, and m=2, but How do you get the answer?

OpenStudy (campbell_st):

differentiate and the sub x = 1/2 to find the gradient if y = ln(x) and when x = 1/2 y = ln(1/2) by substituting to gradient 2 point (1/2, ln (1/2)

OpenStudy (anonymous):

How do you calculate In(1/2)?

OpenStudy (campbell_st):

well using log laws... ln(1/2) = ln(1) - ln(2) and ln(1) = 0 so its just - ln(2)

OpenStudy (anonymous):

f(x) = lnx -> f'(x) = 1/x => f'(1/2) = 1/ (1/2) = 2

OpenStudy (anonymous):

Ok. So how do you end up also with -1?

OpenStudy (campbell_st):

point gradient formula is \[y - y _{1} = m(x - x _{1})\] y - (-ln2)=2(x - 1/2) expand and simplfy y + ln(2) = 2x - 1 y = 2x - ln(2) -1

OpenStudy (anonymous):

Thank you. What is this formula actually used for? I mean, what does the -1 acually represent?

OpenStudy (anonymous):

y = mx + b = 2x + ( ln2 -1)

OpenStudy (campbell_st):

the y intercept is y = -ln2 - 1 ln2 is a number but for convenience its left in its exact form

OpenStudy (anonymous):

I mean 2x + (- ln2 -1)

OpenStudy (anonymous):

Hmm.. So how would you know not to stop at just 2x-In2 and carry on to get -1?

OpenStudy (campbell_st):

its a bit of experience but its also about realising ln2 and 1 are both numbers... so combined they give the y intercept

OpenStudy (anonymous):

Yes. In another question, however, There is no need to do the point-formula, because it stops at just -1... Do In numbers always need the point formula?

OpenStudy (anonymous):

since tangent line is actually straigh line

OpenStudy (anonymous):

Which contact with the curve at specific point

OpenStudy (campbell_st):

yes to find the point for the 2nd tangent... given y = 2ln(2x) so its x = 1/2 and y = 2ln(2 x 1/2) = y = 2ln(1) and ln (1) = 0. the 2nd line passes through the origin

OpenStudy (anonymous):

As long as you apply 1st derivative at specific point --> slope

OpenStudy (anonymous):

Next plug coordinate of contact point --> y intercept b

OpenStudy (anonymous):

Combine slope m and b value => tangent line!

OpenStudy (anonymous):

Thanks! I think I understand it now. I've done two more. Can you help me with one more that I'm posting up. I'm not sure of something?

OpenStudy (anonymous):

Which one?

OpenStudy (anonymous):

I'm now posting

OpenStudy (campbell_st):

ok thats fine

OpenStudy (anonymous):

Thank, campbell :)